# PARCO: Learning Parallel Autoregressive Policies for Efficient Multi-Agent Combinatorial Optimization

### v1

@@ Abstract
Multi-agent combinatorial optimization problems such as routing and scheduling have great practical relevance but present challenges due to their NP-hard combinatorial nature, hard constraints on the number of possible agents, and hard-to-optimize objective functions. This paper introduces PARCO (Parallel AutoRegressive Combinatorial Optimization), a novel approach that learns fast surrogate solvers for multi-agent combinatorial problems with reinforcement learning by employing parallel autoregressive decoding. We propose a model with a Multiple Pointer Mechanism to efficiently decode multiple decisions simultaneously by different agents, enhanced by a Priority-based Conflict Handling scheme. Moreover, we design specialized Communication Layers that enable effective agent collaboration, thus enriching decision-making. We evaluate PARCO in representative multi-agent combinatorial problems in routing and scheduling and demonstrate that our learned solvers offer competitive results against both classical and neural baselines in terms of both solution quality and speed. We make our code openly available at https://github.com/ai4co/parco.
[…]
@@ 1 Introduction
Combinatorial optimization (CO) problems, such as routing and scheduling problems, involve determining an optimal sequence of actions in a combinatorial space and have applications ranging from warehouse operatons (Xie et al., 2023) to network design (Chabarek et al., 2008) and safety-critical systems (Girardey et al., 2010). CO problems are notoriously hard to solve and cannot generally be solved optimally in polynomial time, i.e., they are NP-hard (Jünger et al., 1995). Multi-agent settings gained significant interest due to their applicability in realistic scenarios, such as path finding (Reijnen et al., 2020), drone routing (Ann et al., 2015), disaster management (Bektas, 2006; Cheikhrouhou and Khoufi, 2021) and order delivery (Yakıcı and Karasakal, 2013; Archetti and Bertazzi, 2021), but present even more challenges due to additional constraints and different optimization objectives, including minimizing a global lateness objective or the makespan (Mahmoudinazlou and Kwon, 2024).
While traditional algorithmic methods have significantly contributed to solving a range of problems (Laporte and Osman, 1995; Hejazi and Saghafian, 2005), these approaches often concentrate on single-agent scenarios. With the advent of modern computational techniques, neural network-based approaches have started to yield promising results for complex CO problems in a field known as Neural Combinatorial Optimization (NCO). In particular, Reinforcement Learning (RL) has shown promise due to its ability to learn directly from interactions with environments instead of relying on (costly) labeled datasets in the shape of optimal solutions (Bello et al., 2016; Kwon et al., 2020; Kim et al., 2023a).
Among NCO methods, Autoregressive (AR) sequence generation has gained attention for its ability to manage hard constraints (Kool et al., 2018; Kwon et al., 2021). In the context of NCO, this capability is crucial for addressing complex problems with multiple construction constraints, such as the precedence constraints prevalent in scheduling (Zhang et al., 2020a) as well as pickup and delivery problems (Savelsbergh and Sol, 1995; Parragh et al., 2008). However, one issue of autoregressive sequence generation is the high latency associated with it, especially when considering large problem instances. The problem of high inference latency of AR methods is especially prevalent in the domain of large language models, as multiple stacks of transformer layers have to be computed sequentially for each individual token (Bae et al., 2023).
Figure 1: Solution construction with Autoregressive (top) and Parallel Autoregressive policies (bottom).
Motivated by recent studies on LLMs demonstrating parallel decoding can not only tackle the high generation latency problem of next-token predictors but also improve their answer quality (Qi et al., 2020; Ning et al., 2023; Gloeckle et al., 2024) this paper introduces PARCO (Parallel AutoRegressive Combinatorial Optimization), a novel approach designed to address multi-agent combinatorial problems efficiently. We propose to apply parallel decoding in the NCO domain to increase solution construction efficiency and effectiveness. In contrast to most previous NCO methods, PARCO constructs solutions in parallel for multiple agents, made feasible by a priority-based conflict handler to manage conflicting decisions of different agents. Importantly, we enable the collaborative behavior of agents through specialized inter-agent Communication Layers, essential for finding high-quality solutions to multi-agent combinatorial problems.
Contributions.
We summarize our contributions as follows:
  • We introduce PARCO, a novel approach for tackling multi-agent CO problems by effectively constructing solutions in parallel via a Multiple Pointer Mechanism.
  • We introduce an efficient Priority-based Conflict-Handling scheme to resolve conflict between agents and enhance decision-making.
  • We design a special Communication Layer to help agents communicate and collaborate to obtain a better state-space representation of partial solutions.
  • We show competitive performance against classical solvers and neural baselines in routing and scheduling problems, both in terms of solution quality and computational efficiency.
@@ 2 Related Work
Recent advancements in NCO have shown promising end-to-end solutions for combinatorial optimization problems, as highlighted by Bengio et al. (2021) and Yang and Whinston (2023). NCO has led to the development of aligned neural architectures (Bi et al., 2022; Jin et al., 2023; Luo et al., 2023; Kim et al., 2023c), hybrid methods with OR solvers (Li et al., 2021b; Kim et al., 2024a; Yan and Wu, 2024; Ye et al., 2024a; Kim et al., 2024b), multi-level solution pipelines (Ma et al., 2023a; Li et al., 2023; Xiao et al., 2023; Ye et al., 2023), alongside improved training algorithms (Kim et al., 2023c; Jiang et al., 2023; Drakulic et al., 2023; Sun and Yang, 2023; Gao et al., 2023; Xiao et al., 2023; Li et al., 2024b; Wang et al., 2024) to enhance the heuristic search. These innovations have expanded NCO’s application across a wide array of problems (Chen et al., 2023; Kim et al., 2023b; Zhou et al., 2023; Ma et al., 2023b; Luttmann and Xie, 2024). However, integrating appropriate inductive biases requires manual tuning of model architectures and training algorithms, presenting challenges such as computationally intensive training, the need for specialized hardware, and issues with interpretability and generalizability (Liu et al., 2023).
Multi-agent Constructive Methods for NCO
While several seminal works as Vinyals et al. (2015); Kool et al. (2018); Kwon et al. (2020) propose models that can be used in loose multi-agent settings, such methods cannot be employed directly to model heterogeneous agents with different attributes, such as different capacities or starting locations in routing problems. Zhang et al. (2020b); Falkner and Schmidt-Thieme (2020); Li et al. (2022); Liu et al. (2024c) introduce models to construct solutions of different agents simultaneously. However, the above parallel construction is, in fact, sequential in terms of the model and environment stepping itself, thus not benefitting from true model-side parallel decoding and often lacking in agent collaboration. On the other hand, Son et al. (2024); Zheng et al. (2024) propose to model multi-agent min-max routing as sequential decision-making, allowing for switching the agents after a single-agent solution is complete. Most related to PARCO is Zong et al. (2022), who propose a multi-agent model for the min-sum multi-agent PDP with different decoders. However, this approach is not only tailored to a specific CO problem but also a specific number of agents due to both different decoders for each agent and a non-flexible context embedding, which is dependent on the agent number, while PARCO is more flexible and has a more general communication representation. Moreover, while Zong et al. (2022) manages conflicts by randomly giving precedence to one agent, PARCO learns how to prioritize agents with higher probabilities of selecting a specific node.
3 Background
@@ 3.1 Markov Decision Processes
CO problems can be framed as Markov Decision Processes (MDPs). In this formulation, the problem is defined by a set of states $\mathcal{S}$, where each state $s_{t}\in\mathcal{S}$ represents the configuration of the problem at time step $t$. At each step, an agent selects an action $a_{t}$ from the action space $\mathcal{A}$ according to a policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$, which maps states to actions. The system then transitions from state $s_{t}$ to state $s_{t+1}$ according to a transition function $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$. The agent reaches the terminal state once it has generated a viable solution for the problem instance $\bm{x}$. Typically, a reward $R$ is only obtained in the terminal state and takes the form of the objective function of the respective CO problem.
3.2 Autoregressive Methods for NCO
Given the sequential nature of MDPs, autoregressive (AR) methods pose a natural choice for the policy $\pi$. AR methods construct a viable solution by sequentially generating actions based on the current state and previously selected actions. Formally, the process can be represented as:
$p_{\theta}(\bm{a}|\bm{x})$ $\triangleq\prod_{t=1}^{T}g_{\theta}(a_{t}|a_{t-1},...,a_{0},\bm{h})$ (1)
where $\bm{h}=f_{\theta}(\bm{x})$ is the encoding of problem instance $\bm{x}$ via encoder network $f$, which is used to decode actions autoregressively via decoder $g_{\theta}$. $p_{\theta}$ is a solver that maps $\bm{x}$ to a solution $\bm{a}$. $\bm{a}=(a_{1},...,a_{T})$ represents the (optimal) actions executed in $T$ construction steps, resulting in a feasible solution to CO problems.
Notably, the same AR scheme has been applied to solve multi-agent CO problems, such as vehicle routing problems (Kool et al., 2018; Kwon et al., 2020), in which a single agent tasked with visiting all nodes under specific route length constraints, distinguishing it from a real multi-agent scenario by focusing on optimizing a solitary agent’s path within set limits. To tackle such problems, AR methods sequentially construct multiple routes as if handling a single agent at a time. However, inducing collaborative behavior among agents using the AR method poses a challenge, as it inherently focuses on one agent at a time. This limitation restricts the method’s ability to directly foster interactions among the agents, which is crucial for optimizing collective outcomes in multi-agent scenarios.
3.3 Training AR Models via Reinforcement Learning
In this work, we focus on Reinforcement Learning (RL) solvers for training $p_{\theta}$ as they can be trained without relying on (often hard-to-obtain) labeled solutions.
We can thus cast the optimization problem as follows:
@@ 3.3 Training AR Models via Reinforcement Learning
where $P(\bm{x})$ is problem distribution, $R(\bm{a},\bm{x})$ is reward of $\bm{a}$ given $\bm{x}$. We can solve Eq. 2 by employing various RL methods such as variations of policy gradients (REINFORCE) (Kwon et al., 2020; Kim et al., 2022).
@@ 4 Methodology
4.1 Parallel Multi-Agent Environments
We propose to step multiple actions over the environment simultaneously to enhance efficiency in practice, which can reduce the number of steps required compared to single-agent stepping. In the general multi-agent CO setting, agents $k=1,...,m$ are selecting actions $\bm{a}_{t}=(a_{t}^{1},...,a_{t}^{m})$ from a shared action space $\mathcal{A}$ in parallel at a given decoding step $t$. Given the agent actions $\bm{a}_{t}$, the state of the problem instance transitions from $s_{t}$ to $s_{t+1}$ via some transition function $T:\mathcal{S}\times\mathcal{A}_{1}\times...\times\mathcal{A}_{m}\rightarrow\mathcal{S}$, which usually follows deterministic rules in case of non-stochastic CO problems. After reaching the terminal state, the agents receive a shared reward $R(A,\bm{x})$, with $\bm{x}$ the problem instance and $A=(\bm{a}_{1},...,\bm{a}_{T})$ the sequence of agents actions, which is the objective of the respective CO problem. Fig. 1 illustrates an example of AR and Parallel AR solution construction.
In this work, we propose a model architecture, PARCO, to solve such cooperative multi-agent CO problems efficiently. Drawing on studies on multi-agent RL (MARL) which promotes parameter sharing in cooperative MARL settings with a shared action space (Yu et al., 2022), PARCO establishes a central auto-regressive policy $\pi_{\theta}:\mathcal{S}\rightarrow P(\mathcal{A}_{1}\times...\times\mathcal{A}_{m})$ to decode the next action for the $m$ agents in parallel. This not only makes PARCO agnostic to the number of agents prevalent in the respective CO problem, but also enhances both solution quality and construction speed.
4.2 Parallel Autoregressive Model
In this section, we present our general architecture to solve CO problems efficiently cast as multi-agent sequential decision-making problems. PARCO follows the general encoder-decoder architecture prevalent in autoregressive NCO (Kool et al., 2018; Bello et al., 2016; Vinyals et al., 2015).
Figure 2: Overview of the PARCO model.
Encoder
The job of the encoder in autoregressive NCO approaches is to learn a mapping from the graph representing the CO problem to a higher dimensional embedding space. This typically involves a problem depending on the initial embedding layer, projecting the graph nodes using their feature representations. This is also the case for PARCO, whose encoder first maps features of the graph instance to an embedding space to learn deep relationships between the nodes through a stack of transformer blocks, similar to Kool et al. (2018). The core of each transformer block is a multi-head attention (MHA) layer, which enables message passing between the nodes in the graph and can be defined as follows (Vaswani et al., 2017):
$\text{MHA}(Q,K,V)$ $=\left(\mathop{\big{\|}}_{i=1}^{h}\text{Attn}(QW^{Q}_{i},KW^{K}_{i}VW^{V}_{i})\right)W^{O}$
where
$\text{Attn}(Q,K,V)$ $=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V\,.$

### v2

# Parallel AutoRegressive Models for Multi-Agent Combinatorial Optimization

@@ Abstract
Combinatorial optimization problems involving multiple agents are notoriously challenging due to their NP-hard nature and the necessity for effective agent coordination. Despite advancements in learning-based methods, existing approaches often face critical limitations, including suboptimal agent coordination, poor generalizability, and high computational latency. To address these issues, we propose Parallel AutoRegressive Combinatorial Optimization (PARCO), a reinforcement learning framework designed to construct high-quality solutions for multi-agent combinatorial tasks efficiently. To this end, PARCO integrates three key components: (1) transformer-based communication layers to enable effective agent collaboration during parallel solution construction, (2) a multiple pointer mechanism for low-latency, parallel agent decision-making, and (3) priority-based conflict handlers to resolve decision conflicts via learned priorities. We evaluate PARCO in multi-agent vehicle routing and scheduling problems where our approach outperforms state-of-the-art learning methods and demonstrates strong generalization ability and remarkable computational efficiency.
Machine Learning, ICML
@@ 1 Introduction
Combinatorial optimization (CO) problems involve determining an optimal sequence of actions in discrete spaces with several crucial domains, including logistics and supply chain management (Weinand et al., 2022). Many practical CO problems require a solution to be constructed by coordinating multiple distinct entities (i.e., agents), each with unique characteristics. We call such problems multi-agent CO. This class of problems naturally arises in real-world applications such as coordinated vehicle routing for disaster management (Liberatore et al., 2014), manufacturing (Jungwattanakit et al., 2008) and last-mile delivery optimization (Archetti & Bertazzi, 2021), where heterogeneous agents must operate under complex constraints.
CO problems are notoriously hard to solve and cannot generally be solved optimally in polynomial time due to their NP-hardness (Jünger et al., 1995; Peres & Castelli, 2021). While traditional methods such as exact and heuristic methods have been developed to solve a variety of problems (Laporte & Osman, 1995; Hejazi & Saghafian, 2005), these approaches often concentrate on single-agent scenarios and require long execution times. Moreover, multi-agent CO poses numerous additional challenges such as additional constraints and different optimization objectives including minimizing lateness or the makespan among agents (Soman & Patil, 2020; Mahmoudinazlou & Kwon, 2024).
Recently, Neural Combinatorial Optimization (NCO) has emerged as a powerful alternative by learning efficient neural solvers (Bengio et al., 2021). In particular, Reinforcement Learning (RL) has shown promise due to its ability to learn directly from interactions with CO problems instead of relying on costly labeled datasets in the shape of optimal solutions, and could even outperform traditional approaches by automatically discovering better (neural) heuristics (Bello et al., 2016; Kwon et al., 2020; Kim et al., 2023a). Among NCO methods, Autoregressive (AR) models—constructing solutions step-by-step—have garnered attention for their ability to generate solutions for a variety of problems with hard constraints (Kool et al., 2019a; Kwon et al., 2021; Bi et al., 2024). This capability is crucial for addressing complex problems with multiple constraints, such as heterogeneous capacity (Baldacci et al., 2008) or precedences in pickup and delivery (Parragh et al., 2008) and machine scheduling problems (Zhang et al., 2020a).
However, despite their promising results in single-agent settings, AR-CO methods pose two challenges that hinder their practical adoption in solving complex multi-agent CO problems. First, existing AR methods for multi-agent CO suffer from poor coordination, resulting in unsatisfactory solution quality and poor generalization across varying problem sizes and agent configurations (Zong et al., 2022; Son et al., 2024; Zheng et al., 2024a; Liu et al., 2024b). Moreover, AR sequence generation is associated with high latency due to each single action (or “token”) depending on each single previous one, akin to large language models (Leviathan et al., 2023; Gao et al., 2024). This issue becomes increasingly pronounced when dealing with large problem instances involving numerous agents.
Figure 1: Example multi-agent vehicle routing problem. [Top] Autoregressive policies construct solutions one agent at a time, with inefficient coordination and requiring several steps. [Bottom] Our parallel autoregressive policies ensure not only improved decision-making thanks to inter-agent communication that enables effective coordination but also faster construction in fewer steps.
This paper introduces PARCO (Parallel AutoRegressive Combinatorial Optimization), a novel learning framework to address multi-agent combinatorial problems effectively via parallel solution construction as illustrated in Figure 1. We design specialized transformer-based Communication Layers to enhance coordination among agents during decision-making, enabling collaborative behaviors. Our model leverages a Multiple Pointer Mechanism to efficiently generate actions for various agents simultaneously at each solution construction step. Priority-based Conflict Handlers ensure feasibility and resolve potential conflicts among agents’ decisions based on learned priorities.
We summarize PARCO’s contributions as follows:
  • We propose a general Parallel Autoregressive framework for solution construction in multi-agent CO.
  • We introduce Communication Layers to enhance agent coordination at each parallel construction step.
  • We design a Multiple Pointer Mechanism that reduces latency by efficiently decoding solutions in parallel.
  • We enhance solution quality with Priority-based Conflict Handlers that tie break with learned priorities.
  • We evaluate PARCO on multi-agent vehicle routing and scheduling, where we outperform state-of-the-art learning methods across diverse problems [^1] Code at: https://github.com/ai4co/parco.
@@ 2 Related Work
Recent advancements in Neural Combinatorial Optimization (NCO) have shown promising end-to-end solutions for combinatorial optimization problems (Bengio et al., 2021; Yang & Whinston, 2023). NCO has led to the development of a variety of methods for diverse problems, including the incorporation of problem-specific biases (Jin et al., 2023; Luo et al., 2023; Kim et al., 2023c, b; Son et al., 2023; Gao et al., 2023), bi-level solution pipelines (Li et al., 2021b; Ye et al., 2023b; Zheng et al., 2024b), learning-guided search (Ye et al., 2023a; Kim et al., 2024; Sun & Yang, 2023; Li et al., 2024b, 2025), improvement methods (Hottung et al., 2021; Ma et al., 2021, 2024), effective training algorithms (Kim et al., 2022; Grinsztajn et al., 2023; Drakulic et al., 2023; Corsini et al., 2024; Pirnay & Grimm, 2024; Luo et al., 2024), downstream applications (Chen et al., 2023; Zhou et al., 2023; Ma et al., 2023; Luttmann & Xie, 2024; Yan & Wu, 2024), and the recent development of end-to-end foundation models (Liu et al., 2024a; Zhou et al., 2024; Drakulic et al., 2024; Berto et al., 2024b; Jiang et al., 2024; Li et al., 2024a). Among such methods, RL-based end-to-end autoregressive (AR) models present several advantages, including eliminating the need for labeled solutions, reducing reliance on handcrafted heuristics, and achieving high efficiency in generating high-quality solutions (Kool et al., 2019a; Kwon et al., 2020, 2021; Berto et al., 2024a; Bonnet et al., 2024).
Multi-Agent AR Methods for CO
While the seminal works in AR-CO methods of Vinyals et al. (2015); Kool et al. (2019a); Kwon et al. (2020, 2021) propose models that can be used in loose multi-agent settings, such methods cannot be employed directly to model heterogeneous agents with different attributes and constraints. Building on Kool et al. (2019a), Son et al. (2024) and Zheng et al. (2024a) introduce attention-based policies for multi-agent min-max routing. These models adopt a sequential AR construction strategy, solving for one agent at a time and switching agents only after completing a single-agent solution. While this approach outperforms decentralized methods (Cao et al., 2022; Park et al., 2023) it remains inherently sequential. In contrast, other multi-agent AR methods determine a location for every agent in each solution construction step, but the different agents select their actions sequentially in either random (Zhang et al., 2020b) or learned order (Falkner & Schmidt-Thieme, 2020; Li et al., 2022; Liu et al., 2024b), making these models still suffer from high generation latency as well as missing inter-agent communication and coordination. Zong et al. (2022) propose a multi-agent pickup and delivery model with parallel decoding, using distinct decoders for each agent. However, this approach exhibits limited generalizability due to inflexible fixed decoders for specific agents, lacking a powerful communication mechanism, and conflict resolution handled naively by assigning random precedence to agents, restricting the robustness of the model. PARCO aims at addressing the shortcomings of previous works by leveraging parallel solution construction for any number of agents efficiently with a Multiple Pointer Mechanism, enhancing coordination via Communication Layers, and solving conflicts in a principled manner via Priority-based Conflict handlers. Finally, unlike previous works, PARCO tackles a general framework tackling multi-agent CO without restricting to a single class of problems.
3 Preliminaries
@@ 3.1 Markov Decision Processes
Multi-agent CO problems can be formulated as Markov Decision Processes (MDPs) and solved autoregressively using RL (Mazyavkina et al., 2021). In this framework, a solution $\bm{a}$ to a CO problem instance $\bm{x}$ is represented as a sequence of actions. Actions $a_{t}$ are selected sequentially from the action space $\mathcal{A}$ of size $N$ based on the current state $s_{t}\in\mathcal{S}$, which encodes the problem’s configuration at step $t$. In multi-agent problems, at each step one agent $m\in\mathcal{M}=\{1,\dots,M\}$ selects an action $a^{m}_{t}$ according to a policy $\pi_{\theta}$, usually represented by a $\theta$-parametrized neural network, mapping states to actions. Agents are selected either by some predefined precedence rule as in the sequential planning of Son et al. (2024), where agent solutions are constructed one after another, or by the policy itself, in which case $\pi_{\theta}:\mathcal{S}\rightarrow\mathcal{A}\times\mathcal{M}$. Given the agent and its corresponding action, the problem then transitions from state $s_{t}$ to state $s_{t+1}$ according to a transition function $\tau:\mathcal{S}\times\mathcal{A}\times\mathcal{M}\rightarrow\mathcal{S}$. This process reaches the terminal state once it has generated a feasible solution $\bm{a}=(a_{1},...,a_{T})$ for the problem instance $\bm{x}$ in $T$ construction steps. The (sparse) reward $R(\bm{a},\bm{x})$ is usually obtained only in the terminal state and takes the form of the negative of the cost function of the respective CO problem.
3.2 AR Models for CO
Given the sequential nature of MDPs, AR methods pose a natural choice for the policy $\pi_{\theta}$. AR methods construct a viable solution by sequentially generating actions based on the current state and previously selected actions. Without loss of generality, the process can be represented in an encoder-decoder framework as:
$p_{\theta}(\bm{a}|\bm{x})$ $\triangleq\prod_{t=1}^{T}g_{\theta}(a_{t}|\bm{a}_{<t},\bm{h})$ (1)
where $\bm{a}_{<t}=(a_{1},\ldots,a_{t-1})$ is the sequence of actions taken prior to $t$ and $\bm{h}=f_{\theta}(\bm{x})$ is an encoding of the problem instance $\bm{x}$ obtained via the encoder network $f$. The decoder $g_{\theta}$ then autoregressively generates the sequence of actions, conditioned on $\bm{h}$ and the previously generated actions. The parameters $\theta$ encompass both the encoder and decoder components, which together define the policy as the mechanism for producing the joint distribution $p_{\theta}(\bm{a}|\bm{x})$.
3.3 RL Training
RL leverages the MDP formulation to learn the parameters $\theta$ of the policy $\pi_{\theta}$ without relying on labeled optimal solutions. We can thus cast the optimization problem as follows:
@@ 3.3 RL Training
where $P(\bm{x})$ is the distribution over problem instances and $R(\bm{a},\bm{x})$ is the reward of a solution. We can solve the optimization problem in Equation 2 using various RL methods, including policy gradient (REINFORCE) variations (Kool et al., 2019b; Kwon et al., 2020; Kim et al., 2022).
@@ 4 Methodology
We now outline the general structure of PARCO as shown in Figure 2. First, we formally define parallel multi-agent MDPs for CO (Section 4.1), which we use as a basis to derive our overall parallel autoregressive approach (Section 4.2). We then describe in detail the components of our model: Multi-Agent Encoder (Section 4.3), Communication Layers (Section 4.4), Decoder with Multiple Pointer Mechanism (Section 4.5) and Conflict Handlers (Section 4.6). Finally, we outline the training scheme (Section 4.7).
Figure 2: Overview of PARCO. Our model encodes multi-agent CO problems into separate agent and node embeddings. Communication Layers allow for coordination among agents during decoding, which enhances solution quality. Actions are decoded efficiently autoregressively in parallel through a Multiple Pointer Mechanism enhanced by a Priority-based Conflict Handler.
4.1 Parallel Multi-Agent MDPs
We reformulate the multi-agent MDPs of Section 3.1 as parallel multi-agent MDPs by selecting multiple actions from a joint action space simultaneously to enhance efficiency and coordination (Yu et al., 2022). At each step $t$, ${M}$ agents select actions $\bm{a}_{t}=(a_{t}^{1},...,a_{t}^{M})$ according to a policy $\pi_{\theta}:\mathcal{S}\rightarrow\mathcal{A}_{1}\times...\times\mathcal{A}_{M}$, which maps the state space $\mathcal{S}$ to the joint action space of agents. A conflict handling function $\psi:(\mathcal{A}_{1}\times...\times\mathcal{A}_{M})\rightarrow(\mathcal{A}_{1}\times...\times\mathcal{A}_{M})$ ensures action compatibility by allowing only one agent to execute when multiple agents select mutually conflicting actions (e.g., the same customer location), and assigning a fallback action (e.g., staying at current position) to the others. Given the resolved agent actions $\tilde{\bm{a}}_{t}=\psi(\bm{a}_{t})$, the state of the problem $s_{t}$ progresses to $s_{t+1}$ according to the transition function $\tau:\mathcal{S}\times(\mathcal{A}_{1}\times...\times\mathcal{A}_{M})\rightarrow\mathcal{S}$. The agents receive a shared reward $R(\bm{a},\bm{x})$, with $\bm{a}=(\tilde{\bm{a}}_{1},...,\tilde{\bm{a}}_{T})$ the sequence of joint agent actions.
In the following, we refer to the entities corresponding to actions in the MDP formulation of a CO problem (e.g., customer locations in VRPs) as nodes, following the convention of Kool et al. (2019a) and Kwon et al. (2020).
4.2 Parallel AR Models for CO
Motivated by the nature of multi-agent MDPs, PARCO introduces a Parallel AR model for the policy $\pi_{\theta}$. PARCO constructs feasible solutions by simultaneously generating multiple agent actions based on the current state. We formulate the solution generation process in an encoder-decoder framework similarly to Equation 1:
$p_{\theta}(\bm{a}|\bm{x})$ $\triangleq\prod_{t=1}^{T}\psi\left(\prod_{m=1}^{M}g_{\theta}(a_{t}^{m}|\bm{a}_{<t},\bm{h})\right)$ (3)
where $\bm{h}=f_{\theta}(\bm{x})$ is the encoding of problem instance $\bm{x}$ via encoder network $f_{\theta}$, and decoder $g_{\theta}$ implements the policy $\pi_{\theta}$ to autoregressively generate actions efficiently for all agents in parallel. At each step $t$, the decoder outputs joint actions $\bm{a}_{t}=(a_{t}^{1},...,a_{t}^{M})$ for all $M$ agents with $a_{t}^{m}$ being the sampled action of agent $m$. The conflict resolution function $\psi$ ensures action compatibility by allowing only one agent to execute when multiple agents select mutually conflicting actions, resulting in resolved actions $\tilde{\bm{a}}_{t}=\psi(\bm{a}_{t})$. $p_{\theta}$ is thus a solver that maps $\bm{x}$ to a solution $\bm{a}=(\tilde{\bm{a}}_{1},...,\tilde{\bm{a}}_{T})$.
A benefit of our parallel formulation is that the total number of construction steps $T$ can be substantially lower compared to purely AR methods. While the latter require $\sum_{m=1}^{M}T_{m}$ total actions to construct a solution with $T_{m}$ being the number of steps required by agent $m$ to finish its task, PARCO’s Parallel AR needs only $\max_{m}T_{m}$ steps as agents effectively divide the solution space and act concurrently. This leads to a faster solution construction as illustrated in Figures 1 and 5 and significantly reduced training times.
4.3 Multi-Agent Encoder
The multi-agent encoder $f_{\theta}$ transforms an input instance $\bm{x}$ into a hidden representation $\bm{h}$. In PARCO, we explicitly model agents and employ separate agent and node embedding layers similar to Son et al. (2024) to project agents and nodes into the same embedding space.
The agent embedding layer projects $k_{a}$ agent features – such as vehicle locations and capacities (routing) or machine characteristics (scheduling) – into a $d$-dimensional space using a linear projection ${\mathbf{W}}_{a}\in\mathbb{R}^{k_{a}\times d}$. Let $\bm{h}_{a}^{(0)}=\bm{x}_{a}{\mathbf{W}}_{a}$ denote the initial agent embeddings, where $\bm{x}_{a}\in\mathbb{R}^{M\times k_{a}}$ denotes the matrix of agent features. Similarly, the node embedding layer projects $k_{n}$ node features – such as customer demands in vehicle routing or job durations in scheduling problems – into the same $d$-dimensional space using a linear projection ${\mathbf{W}}_{n}\in\mathbb{R}^{k_{n}\times d}$. The initial embeddings of the nodes are defined as $\bm{h}_{n}^{(0)}=\bm{x}_{n}{\mathbf{W}}_{n}$ where $\bm{x}_{n_{i}}\in\mathbb{R}^{N\times k_{n}}$ represents the node feature matrix.
Depending on the problem structure, the initial agent and node embeddings might either be concatenated as $\bm{h}^{(0)}=\text{Concat}(\bm{h}_{a}^{(0)},\bm{h}_{n}^{(0)})$ and passed through $L$ transformer blocks, consisting of multi-head self-attention $\text{MHA}(\bm{h}^{(0)},\bm{h}^{(0)},\bm{h}^{(0)})$ and multi-layer perceptrons (MLPs) as defined in Vaswani et al. (2017). Or, agent and node embeddings are used separately as query and keys/values, respectively, in a cross-attention mechanism $\text{MHA}(\bm{h}_{a}^{(0)},\bm{h}_{n}^{(0)},\bm{h}_{n}^{(0)})$ akin to MatNet (Kwon et al., 2021). The final embeddings $\bm{h}=\{\bm{h}_{a},\bm{h}_{n}\}$ emitted by the last encoder layer contain processed representations of both agents and nodes that capture their interactions as well as the overall problem structure.
4.4 Communication Layers
At each step $t$ of the decoding process of Equation 3, given the encoded representations $\bm{h}$, we construct dynamic agent queries that capture the current state of both agents and the environment. For each agent $m$, we form a context embedding $\bm{d}_{m}=\text{Concat}(\bm{h}_{a^{m}},\bm{h}_{\delta^{m}_{t}},\bm{h}_{e})$ consisting of the following components: (1) the (static) embedding of the agent $\bm{h}_{a^{m}}$; (2) a projection $\bm{h}_{\delta^{m}_{t}}=\delta^{m}_{t}{\mathbf{W}}_{\delta}\in\mathbb{R}^{d}$ of the agent’s dynamics $\delta^{m}_{t}$ like its current location and capacity (routing) or the time until the agent becomes idle (scheduling); (3) a projection $\bm{h}_{e}=e_{t}{\mathbf{W}}_{\delta}\in\mathbb{R}^{d}$ of the dynamic environment features $e_{t}$ that encode the current problem’s state. These dynamic embeddings are then projected into query vectors $\bm{q}_{m}=\bm{d}_{m}{\mathbf{W}}_{q}$ where ${\mathbf{W}}_{q}\in\mathbb{R}^{3d\times d}$ is a learnable projection matrix.
The resulting queries $\bm{q}=[\bm{q}_{1},\dots,\bm{q}_{M}]\in\mathbb{R}^{M\times d}$ are then processed through communication layers comprising multi-head self-attention followed by an MLP:
$\bm{q}^{\prime}$ $=\text{Norm}(\text{MHA}(\bm{q},\bm{q},\bm{q})+\bm{q})$ (4)
$\bm{q}$ $=\text{Norm}(\text{MLP}((\bm{q}^{\prime}))+\bm{q}^{\prime})$ (5)
where Norm denotes a normalization layer (Ioffe & Szegedy, 2015; Zhang & Sennrich, 2019). These layers are inherently agent-count agnostic, which allows PARCO to handle an arbitrary number of agents, making it more flexible and generalizable across different problems. Communication Layers allow agents to coordinate their actions by attending to both the problem structure and other agents’ states while maintaining efficiency through parallel processing.
4.5 Decoder with Multiple Pointer Mechanism
PARCO’s decoder extends the attention-based pointer mechanism of Kool et al. (2019a) – originally designed for single-agent scenarios and recently applied to sequential multi-agent planning with a single agent at a time (Son et al., 2024; Zheng et al., 2024a) – to handle multiple agents operating in parallel via a Multiple Pointer Mechanism.
Starting with the processed agent queries $\bm{q}$ that underwent communication, we first compute agent-specific representations through masked cross MHA: